# RAGFormer: Learning Semantic Attributes and Topological Structure for Fraud Detection

### v2

@@ 1 Introduction
  • Extensive experiments on two public opinion fraud datasets and an industrial financial fraud dataset demonstrate the superiority of our method. RAGFormer shows substantial improvement on large-scale datasets especially, with up to 11.59% improvement over state-of-the-art methods.
@@ 5.2 Quantitative Results
Table 2 and Table 3 present the overall performance of the proposed method RAGFormer and other baselines on opinion and financial fraud detection respectively. Table 2 shows that RAGFormer achieves the state-of-the-art (SOTA) performance compared to previous methods on the YelpChi dataset, which better reflects the performance gaps between algorithms. Specifically, RAGFormer surpasses the previous best method GAGA, with 11.59% AP improvement, which demonstrates the effectiveness of merging semantic and topological features from a holistic view. Moreover, on the relatively smaller dataset, Amazon, RAGFormer still maintains the lead on all metrics compared to most methods and performs comparably with COFRAUD [39], a strong baseline that focuses on encoding multi-relation interactions. On the financial fraud detection dataset Pay, RAGFormer again shows its superiority on all metrics on Table 3, where RAGFormer outperforms GAGA with $\approx$ 3%-5% across all metrics.
@@ 5.3 Ablation Study
To validate how each module of RAGFormer benefits the fraud detection performance, we conducted an ablation experiment to evaluate different components of RAGFormer. We present three baselines here: 1) GAGA, which is the Transformer-based state-of-the-art method; 2) Semantic Encoder, which removes the topology encoder from RAGFormer; and 3) Topology Encoder, which removes the semantic encoder from RAGFormer. We experiment with all three baselines and RAGFormer on YelpChi and Pay datasets which possess more complex graphs than Amazon and amplify the differences between different methods. As shown in Table 4, each component contributes to the proposed method. Semantic Encoder is slightly better than GAGA in all metrics on both datasets, which illustrates that gradually aggregating multi-relation information is beneficial for learning the semantic features. Using the topology encoder alone (Relation-Aware GCN) leads to a performance drop by 7.4% in AUC, 27.5% in AP, and 17.2% in F1-macro on the YelpChi dataset, while it outperforms the semantic encoder on the larger dataset Pay in all metrics. This indicates that the relative importance of semantic features and topological features varies with the data. Among them, the topological features are more suitable in larger datasets, where the relation between nodes is more complex. Moreover, the advocated RAGFormer, which integrates the topology encoder and the semantic encoder, always harvests the best performance in all metrics across both datasets. These results further emphasize the need to focus on both aspects of the graph. Our Relation-Aware GNN architecture can holistically capture topological information, making it an effective supplement to the semantic encoder, regardless of the size of the dataset.
@@
There are a variety of effective methods to merge features from two embedding spaces. Therefore, in this ablation study, we verify whether self-attention with residual connection is the first choice for the fusion of semantic and topological features in fraud detection task. We consider four other feature fusion methods: (1) “Concat”: directly concatenating the two features; (2) “Add”: directly adding two feature vectors; (3) “Gated”: using a learnable gate unit like GRU [2]. We use a linear layer followed by a sigmoid function to learn the weight $\gamma\in\mathbb{R}^{d}$ from $\boldsymbol{X}_{sem}$. In this setting, the final embedding $\boldsymbol{X}_{final}$ can be formulate as: $\boldsymbol{X}_{final}=\gamma*\boldsymbol{X}_{sem}+(1-\gamma)*\boldsymbol{X}_{gcn}$ , where $\boldsymbol{X}_{sem}$ represents the semantic features and $\boldsymbol{X}_{gcn}$ represents the topological features. 4) “Attention w/ res”: a variant of our fusion module that removes the residual connection. Similar to previous experiments, we use the same setting and apply the five methods on the YelpChi and Pay datasets. We present the experiment results in Table 5. It is evident that the direct concatenation and gate unit fall behind the self-attention but with a relatively smaller gap than simple addition. However, these two methods lack the cross-modality handling ability that attention-based mechanisms have, thus leading to a suitable application of self-attention for fusing two features from different embedding spaces. Furthermore, our experiments demonstrate that ’Attention w/ res’ outperforms ’Attention w/o res’ across all metrics, highlighting the effectiveness of residual connection in ensuring more robust fusion. In Addition, the residual connection can also accelerate the convergence of RAGFormer. We present more ablation experiment results in Appendix A.4
@@ 5.5 Parameter Analysis
Training Ratio Analysis. In a semi-supervised setting, the proportion of the training set significantly influences the model’s performance. To assess the robustness of the proposed algorithm in this context, we conducted experiments by varying the proportion of the training set. Following previous works [3, 15, 33], we set the proportions of the training set to 10%, 20%, 30%, 40% respectively. We compare the results of RAGFormer against GAGA and CAREGNN on the YelpChi dataset since CAREGNN is a widely used fraud detector with open-source code. As depicted in Figure. 6, RAGFormer exhibits remarkable performance across various training ratios, outperforming other methods by a substantial margin. In every setting, RAGFormer achieves the highest results. Notably, our method attains performance comparable to that of GAGA using merely 10% of the training data, whereas GAGA requires 40%. This verifies that with the fusion of semantic and topological features, RAGFormer can make accurate judgments on node’s belongings with much less data, which is highly advantageous for application on large graphs or in scenarios with limited annotated data. Furthermore, our approach attains its best performance at 30% training ratio, indicating its remarkable data utilization capability. More results on Pay are provided in Appendix 6.
@@ A.4 Ablation on GNN Backbone
A.5 Hyper-parameter Sensitivity Analysis
We present the full results of the hyper-parameter sensitivity Analysis, including results on Pay.
@@ A.5 Hyper-parameter Sensitivity Analysis
Training Ratio Analysis. In a semi-supervised setting, the proportion of the training set significantly influences the model’s performance. To assess the robustness of the proposed algorithm in this context, we conducted experiments by varying the proportion of the training set. Following previous works [3, 15, 33], we set the proportions of the training set to 10%, 20%, 30%, 40% respectively. We compare the results of RAGFormer against GAGA and CAREGNN on YelpChi dataset since CAREGNN is a widely used fraud detector with open-source code. As depicted in Figure 6, RAGFormer exhibits remarkable performance across various training ratios, outperforming other methods by a substantial margin. In every setting, RAGFormer achieves the highest results in all metrics. Notably, our method attains performance comparable to that of GAGA using merely 10% of the training data, whereas GAGA requires 40%. This verifies that with the fusion of semantic and topological features, RAGFormer can make accurate judgments on node’s belongings with much less data, which is highly advantageous for application on large graphs or in scenarios with limited annotated data. Furthermore, our approach attains its best performance at 30% training ratio, indicating its remarkable data utilization capability.
On the Pay dataset, RAGFormer performs best with only 20% training data across the other two baselines using double training data, which also indicates that RAGFormer is not hungry for data and converges fast.

### v3

@@ 1 Introduction
  • Extensive experiments on two public opinion fraud datasets and an industrial financial fraud dataset demonstrate the superiority of our method. RAGFormer shows substantial improvement on large-scale datasets especially, with up to 12.08% improvement over state-of-the-art methods.
@@ 5.2 Quantitative Results
Table 2 and Table 3 present the overall performance of the proposed method RAGFormer and other baselines on opinion and financial fraud detection respectively. Table 2 shows that RAGFormer achieves the state-of-the-art (SOTA) performance compared to previous methods on the YelpChi dataset, which better reflects the performance gaps between algorithms. Specifically, RAGFormer surpasses the previous best method GAGA, with 12.08% AP improvement, which demonstrates the effectiveness of merging semantic and topological features from a holistic view. Moreover, on the relatively smaller dataset, Amazon, RAGFormer still maintains the lead on all metrics compared to most methods and performs comparably with COFRAUD [39], a strong baseline that focuses on encoding multi-relation interactions. On the financial fraud detection dataset Pay, RAGFormer again shows its superiority on all metrics on Table 3, where RAGFormer outperforms GAGA with $\approx$ 3%-5% across all metrics.
@@ 5.3 Ablation Study
To validate how each module of RAGFormer benefits the fraud detection performance, we conducted an ablation experiment to evaluate different components of RAGFormer. We present three baselines here: 1) GAGA, which is the Transformer-based state-of-the-art method; 2) Semantic Encoder, which removes the topology encoder from RAGFormer; and 3) Topology Encoder, which removes the semantic encoder from RAGFormer. We experiment with all three baselines and RAGFormer on YelpChi and Pay datasets which possess more complex graphs than Amazon and amplify the differences between different methods. As shown in Table 4, each component contributes to the proposed method. Semantic Encoder is slightly better than GAGA in all metrics on both datasets, which illustrates that gradually aggregating multi-relation information is beneficial for learning the semantic features. Using the topology encoder alone (Relation-Aware GCN) leads to a performance drop by 7.5% in AUC, 28.0% in AP, and 17.2% in F1-macro on the YelpChi dataset, while it outperforms the semantic encoder on the larger dataset Pay in all metrics. This indicates that the relative importance of semantic features and topological features varies with the data. Among them, the topological features are more suitable in larger datasets, where the relation between nodes is more complex. Moreover, the advocated RAGFormer, which integrates the topology encoder and the semantic encoder, always harvests the best performance in all metrics across both datasets. These results further emphasize the need to focus on both aspects of the graph. Our Relation-Aware GNN architecture can holistically capture topological information, making it an effective supplement to the semantic encoder, regardless of the size of the dataset.
@@
There are a variety of effective methods to merge features from two embedding spaces. Therefore, in this ablation study, we verify whether self-attention with residual connection is the first choice for the fusion of semantic and topological features in fraud detection task. We consider four other feature fusion methods: (1) “Concat”: directly concatenating the two features; (2) “Add”: directly adding two feature vectors; (3) “Gated”: using a learnable gate unit like GRU [2]. We use a linear layer followed by a sigmoid function to learn the weight $\gamma\in\mathbb{R}^{d}$ from $\boldsymbol{X}_{sem}$. In this setting, the final embedding $\boldsymbol{X}_{final}$ can be formulate as: $\boldsymbol{X}_{final}=\gamma*\boldsymbol{X}_{sem}+(1-\gamma)*\boldsymbol{X}_{gcn}$ , where $\boldsymbol{X}_{sem}$ represents the semantic features and $\boldsymbol{X}_{gcn}$ represents the topological features. (4) “Attention w/ res”: a variant of our fusion module that removes the residual connection. Similar to previous experiments, we use the same setting and apply the five methods on the YelpChi and Pay datasets. We present the experiment results in Table 5. It is evident that the direct concatenation and gate unit fall behind the self-attention but with a relatively smaller gap than simple addition. However, these two methods lack the cross-modality handling ability that attention-based mechanisms have, thus leading to a suitable application of self-attention for fusing two features from different embedding spaces. Furthermore, our experiments demonstrate that ’Attention w/ res’ outperforms ’Attention w/o res’ across all metrics, highlighting the effectiveness of residual connection in ensuring more robust fusion. In Addition, the residual connection can also accelerate the convergence of RAGFormer. We present more ablation experiment results in Appendix A.4 and A.5
@@ 5.5 Parameter Analysis
Training Ratio Analysis. In a semi-supervised setting, the proportion of the training set significantly influences the model’s performance. To assess the robustness of the proposed algorithm in this context, we conducted experiments by varying the proportion of the training set. Following previous works [3, 15, 33], we set the proportions of the training set to 10%, 20%, 30%, 40% respectively. We compare the results of RAGFormer against GAGA and CAREGNN on the YelpChi dataset since CAREGNN is a widely used fraud detector with open-source code. As depicted in Figure. 6, RAGFormer exhibits remarkable performance across various training ratios, outperforming other methods by a substantial margin. In every setting, RAGFormer achieves the highest results. Notably, our method attains performance comparable to that of GAGA using merely 10% of the training data, whereas GAGA requires 40%. This verifies that with the fusion of semantic and topological features, RAGFormer can make accurate judgments on node’s belongings with much less data, which is highly advantageous for application on large graphs or in scenarios with limited annotated data. Furthermore, our approach attains its best performance at 30% training ratio, indicating its remarkable data utilization capability. More results on Pay are provided in Appendix A.6.
@@ A.4 Ablation on GNN Backbone
A.5 Ablation on the contribution of each relation graph
Table 9: Ablation study on relation graph.
Relation YelpChi
AUC AP F1
R-U-R 0.8740 0.5734 0.7356
R-T-R 0.9123 0.6871 0.7586
R-S-R 0.9585 0.8394 0.8358
R-U-R&R-T-R 0.9400 0.7848 0.8307
R-U-R&R-S-R 0.9738 0.8970 0.8993
R-T-R&R-S-R 0.9685 0.8796 0.8798
All 0.9781 0.9222 0.9098
Abaltion on relation graph
Fraud detection methods focus on extracting information from multi-relation graphs. To assess the validity of each relation graph, we conducted an ablation experiment on the YelpChi dataset to determine the contribution of each individual relation. The YelpChi dataset includes three types of relations, resulting in a total of six subgraphs: three single-relation graphs and three dual-relation graphs. We applied RAGFormer on all six subgraphs as part of our ablation experiments. When only one relationship is present, the Relation-Aware GCN degenerates into a single GCN. The results of RAGFormer on these subgraphs are shown in Table 9. Our findings are as follows: (1) The results indicate that utilizing a single relation or a pair of relations yields inferior performance compared to leveraging all three relations simultaneously. Furthermore, using two relations concurrently outperforms using a single relation. This demonstrates that the three relations effectively complement one another and that incorporating all three relations is the most efficient approach. (2) Each time an additional relation is incorporated, the performance shows a notable improvement. This indicates that our RAGFormer can effectively leverage multiple relations, making it well-suited for learning on complex multi-relation graphs. (3) Even with only one relation (R-S-R), our RAGFormer can significantly outperform the state-of-the-art method GAGA, which fully proves the advantage of learning both semantic features and topological features. (4) Among the three relations, the R-S-R relation is the most effective. Combining R-S-R with another relation can achieve performance nearly equivalent to those obtained by using all three relations.
@@ A.5 Ablation on the contribution of each relation graph
A.6 Training Ratio Analysis
We present the full results of the training ratio analysis, including results on Pay.
In a semi-supervised setting, the proportion of the training set significantly influences the model’s performance. To assess the robustness of the proposed algorithm in this context, we conducted experiments by varying the proportion of the training set. Following previous works [3, 15, 33], we set the proportions of the training set to 10%, 20%, 30%, 40% respectively. We compare the results of RAGFormer against GAGA and CAREGNN on YelpChi dataset since CAREGNN is a widely used fraud detector with open-source code. As depicted in Figure 6, RAGFormer exhibits remarkable performance across various training ratios, outperforming other methods by a substantial margin. In every setting, RAGFormer achieves the highest results in all metrics. Notably, our method attains performance comparable to that of GAGA using merely 10% of the training data, whereas GAGA requires 40%. This verifies that with the fusion of semantic and topological features, RAGFormer can make accurate judgments on node’s belongings with much less data, which is highly advantageous for application on large graphs or in scenarios with limited annotated data. Furthermore, our approach attains its best performance at 30% training ratio, indicating its remarkable data utilization capability. On the Pay dataset, RAGFormer performs best with only 20% training data across the other two baselines using double training data, which also indicates that RAGFormer is not hungry for data and converges fast.